### v1

# Joint semi-supervised and contrastive learning enables zero-shot domain-adaptation and multi-domain segmentation

Abstract
Despite their effectiveness, current deep learning models face challenges with images coming from different domains with varying appearance and content. We introduce SegCLR, a versatile framework designed to segment volumetric images across different domains, employing supervised and contrastive learning simultaneously to effectively learn from both labeled and unlabeled data. We demonstrate the superior performance of SegCLR through a comprehensive evaluation involving three diverse clinical datasets of retinal fluid segmentation in 3D Optical Coherence Tomography (OCT), various network configurations, and verification across 10 different network initializations. In an unsupervised domain adaptation context, SegCLR achieves results on par with a supervised upper-bound model trained on the intended target domain. Notably, we discover that the segmentation performance of SegCLR framework is marginally impacted by the abundance of unlabeled data from the target domain, thereby we also propose an effective zero-shot domain adaptation extension of SegCLR, eliminating the need for any target domain information. This shows that our proposed addition of contrastive loss in standard supervised training for segmentation leads to superior models, inherently more generalizable to both in- and out-of-domain test data. We additionally propose a pragmatic solution for SegCLR deployment in realistic scenarios with multiple domains containing labeled data. Accordingly, our framework pushes the boundaries of deep-learning based segmentation in multi-domain applications, regardless of data availability —– labeled, unlabeled, or nonexistent.
**footnotetext: Corresponding author: alvaro.gomariz@roche.com
I Introduction
Routine assessment of eye conditions in clinics is being transformed by the applications of deep learning in ophthalmology. These methods facilitate quantitative analysis of 3D images of eyes acquired with Optical Coherence Tomography (OCT) devices [1], which create volumetric datasets by stacking 2D slices known as B-scans. The use of supervised learning in segmentation of retinal fluids, mainly via UNet [2] based neural networks, has led to advances in diagnosis, prognosis, and a deeper understanding of eye diseases, such as neovascular age-related macular degeneration (nAMD) and Diabetic Macular Edema (DME) [3, 4, 5, 6]. However, training such supervised deep neural networks requires large amounts of labeled data, the procurement of which is costly and sometimes infeasible. Most importantly, such labeling needs to be repeated for each problem setting/domain; since trained models often fail when inference data differs from labeled examples, so-called domain-shift, e.g., between images from different eye diseases and different OCT devices [7].
Domain adaptation techniques aim to maintain a similar performance on images from new domains while reducing the annotation effort required in such newer datasets. Unsupervised domain adaptation aims to leverage information learned from a labeled data domain for applications in other domains where only unlabeled data is available. To this end, many deep learning methods have been proposed [8], mostly using generative adversarial networks, e.g., to translate visual appearance across OCT devices [9]. Ultimately, it is desirable for learned models to implicitly generalize to other domains without requiring any new data, i.e., ideally not even unlabeled data from a new domain. This scenario, known as zero-shot domain adaptation [10], provides huge advantages in real-world deployment by obviating any retraining or customization being required per test domain encountered.
Advances in self-supervised learning have been increasingly successful in extracting informative features from images without accessing their labels, hence enabling the exploitation of much larger unlabeled datasets [11, 12, 13, 14, 15, 16, 17, 18]. These methods usually employ different versions of Siamese networks [19] that can be applied to different inputs in parallel while sharing their weights. For self-supervision, typically different augmentations of an image are treated as positive samples, which are desired to map close-by or to a single point on a learned manifold, and therefore minimizing the feature distances between these positive samples when passed through a Siamese network. To prevent such optimization from collapsing to a constant output, different approaches have been proposed: For instance, SimSiam [14] employs a stop-gradient operation for one of the Siamese encoders. Another successful strategy, contrastive learning (CL), introduces the concept of negative pairs, the distances of which should meanwhile be maximized. SimCLR [12] is one of the most widely-adopted CL frameworks. Self-supervised learning is commonly used for model pretraining, typically based on natural images such as ImageNet [20]. Such pretrained model is subsequently fine-tuned or distilled for downstream tasks such as classification, detection, or segmentation [13].
Models pretrained with natural images are of limited use for medical applications, which often involve images with volumetric content and with substantially differing appearances from natural images. This has led to recent application-specific approaches for CL pair generation in medical context [21, 22]. USCL [22] minimizes the feature distance between frames of the same ultrasound video, while maximizing the distance between frames of different videos, in order to produce pretrained models for ultrasound applications. USCL also proposes a joint semi-supervised approach, which simultaneously minimizes a contrastive and supervised classification loss. However, to be applicable for image segmentation, this method relies on subsequent fine-tuning, which is potentially sub-optimal for preserving the unlabeled information for the intended task of segmentation. In fact, there exist little work on CL methods on image segmentation without fine-tuning.
We propose SegCLR as a CL solution to improve segmentation quality by leveraging both manual annotations and unlabeled data. To that end we introduce a semi-supervised framework for joint training of CL together with segmentation labels. Preliminary results of this work have been presented as a conference contribution [23] for two datasets acquired with different OCT scanners, where the domain shift originates from a change in acquisition device. SegCLR without any target-domain labels is shown to achieve unsupervised domain adaptation results close to an upper bound that uses supervised learning on the target domain. Our additional contributions in this work are: (1) In addition to acquisition device change, we present results with a third dataset with a new disease indication, which helps demonstrate the benefits of SegCLR also for domain shift in disease type. (2) An analysis of the amount of unlabeled target data is presented, indicating its limited effect on the resulting segmentation performance. (3) Exploiting the finding above, we propose a zero-shot domain adaptation use-case of our SegCLR framework, for which an extensive analysis reveals substantial benefits showing that SegCLR trained models are implicitly better at generalizing to unseen domains. (4) An analysis of the stability of the results upon distinct network initialisations reveals that most proposed SegCLR variants perform well and similarly, confirming the stability of the overall SegCLR framework to method setting choices such as pair generation and contrastive projection strategies. This further highlights the importance of analyzing many randomized result replicates in assessing deep learning models in general. (5) In a multi-domain training scenario of SegCLR, we show that with labeled data available from multiple domains, SegCLR applied to the domains simultaneously yields results superior to conventional supervised training as well as SegCLR trained on individual domains, potentially better leveraging cross-domain information.
II Methods
The considered learning frameworks are illustrated in Fig. 1. Let a source domain ($D^{s}$) contain labeled data, with which Baseline models can be trained following typical supervised learning schemes. Distinct domains for which only unlabeled images can be observed are considered target domains ($D^{t}$). We first study an unsupervised domain adaptation setting as in [23], where SegCLR is trained for a source domain $D^{s}$ and a target domain $D^{t}$ by using a combination of supervision and contrastive losses. We then investigate its zero-shot domain adaptation application, where SegCLR is trained using only data from a labeled source domain $D^{s}$, while being applied on unseen target domains $D^{t}$, i.e., on domains for which no data (labeled or unlabeled) is used in training.
Figure 1: Illustration of Unsupervised Domain Adaptation (UDA) and Zero-Shot Domain Adaptation frameworks studied herein for SegCLR. The colored ellipses indicate the losses to use in training, with the variable superscripts representing the domain being source (s) or target (t).
II-A Supervised learning
For the segmentation backbone, we adopted the proven UNet architecture [2] (details in Fig. 2), which can be modeled as $F(\cdot)$ processing an image $x$ to predict a segmentation map $p=F(x)$ to approximate an (expert-annotated) ground truth segmentation $y$. $F$ is learned by minimizing a supervised loss $\mathcal{L}_{\mathrm{sup}}$, which in our work is the logarithmic Dice loss of labeled data in a source domain $D^{s}$:
$\mathcal{L}_{\mathrm{sup}}=-\sum_{(p_{i},y_{i})\in D^{s}}\log{\frac{2\sum_{j\in\mathrm{pixels}}y_{i}^{j}p_{i}^{j}}{\epsilon+\sum_{j\in\mathrm{pixels}}(y_{i}^{j}+p_{i}^{j})}}$ (1)
for all training images $(x_{i},y_{i})\in D^{s}$, where $\epsilon$ is a small number to avoid division by 0.
Figure 2: SegCLR architecture employed for joint supervised and contrastive learning. Layers are represented as arrows and their outputs as rectangles. The width and height of these outputs is annotated at the upper left of the rectangles, and the number of features at the bottom. $F(\cdot)$ is the segmentation backbone, $E(\cdot)$ the encoder, and $C(\cdot)$ the contrastive projection. The architectures of $\rho^{\mathrm{agg}}$ and $\rho^{\mathrm{MLP}}$ are described in Section II-E.
II-B Self-supervised learning
Self-supervised learning aims to optimize an encoder $E(\cdot)$ to achieve good representations $h=E(x)$ without the need of manually annotated labels $y$. In the literature (for mostly classification tasks), the ResNet architecture is commonly used for $E(\cdot)$. Instead, we use the existing UNet encoder in order to adapt the learned features $h$ for our intended segmentation task. A subsequent contrastive projection head $C(\cdot)$ then maps the bottleneck-layer features to vector projections $z=C(h)$ on which the contrastive loss $\mathcal{L}_{\mathrm{con}}$ is applied. The architecture used for $E(\cdot)$ and $C(\cdot)$ is illustrated in Fig. 2. We herein employ two widely adopted self-supervised learning frameworks: SimCLR [12] and SimSiam [14].
In SimCLR, $\mathcal{L}_{\mathrm{con}}$ aims to minimize the distance between positive pairs of images and maximize the distance between negative pairs. The positives $(x_{i}^{\prime},x_{i}^{\prime\prime})$ are created from each image $x_{i}$ by a defined pair generator $P(\cdot)$ described further in Section II-D below, i.e. $P(x_{i})\rightarrow(x_{i}^{\prime},x_{i}^{\prime\prime})$. The negatives pairs $(x_{i},x_{k})$ can be formed using other images $x_{k}$ for $k\neq i$. We employ a version of the normalized temperature-scaled cross entropy loss [18] adapted to our problem setting as:
$L_{\mathrm{con}}^{\mathrm{CLR}}=\sum_{P(x_{i}),\ x_{i}\in D}\big{(}\,l(z_{i}^{\prime},z_{i}^{\prime\prime})+l(z_{i}^{\prime\prime},z_{i}^{\prime})\,\big{)}$ (2)
$l(z_{i}^{\prime},z_{i}^{\prime\prime})=-\log\frac{\exp{\big{(}d(z_{i}^{\prime},z_{i}^{\prime\prime})}/\tau\big{)}}{\sum_{x_{k}\in D}\mathbbm{1}_{[k\neq i]}\exp{\big{(}d(z_{i}^{\prime},z_{k})/\tau\big{)}}}$ (3)
where $d(u,v)=(u\cdot v)/(||u||_{2}\,||v||_{2})$ and $\tau$ is the temperature scaling parameter.
In SimSiam, a learnable predictor $Q(\cdot)$ is applied on the projection from one network path to predict that from the other:
$L_{\mathrm{con}}^{\mathrm{Siam}}=-\sum_{x_{i}\in D}\Big{(}d\big{(}Q(z_{i}^{\prime}),z_{i}^{\prime\prime}\big{)}+d\big{(}Q(z_{i}^{\prime\prime}),z_{i}^{\prime}\big{)}\Big{)}$ (4)
where the gradients from the second path are omitted for network weight updates (stopgrad) in back-propagation.
II-C Joint semi-supervised and contrastive learning
We adapt the USCL joint training strategy, which was proposed for US video classification, to our segmentation task on 3D images by combining $\mathcal{L}_{\mathrm{sup}}$ and $\mathcal{L}_{\mathrm{con}}$ in a semi-supervised framework illustrated in Fig. 3. Considering a source domain $D^{\mathrm{s}}$ and a target domain $D^{\mathrm{t}}$, total loss $\mathcal{L}$ is calculated as follows:
$\mathcal{L}=\frac{1}{2}\left(\underset{x\in D^{\mathrm{s}}}{\mathcal{L}_{\mathrm{con}}}+\underset{x\in D^{\mathrm{t}}}{\mathcal{L}_{\mathrm{con}}}\right)+\lambda\underset{(x,y)\in D^{\mathrm{s}}}{\mathcal{L}_{\mathrm{sup}}}$ (5)
where $\lambda$ is a hyperparameter that controls relative contributions.
Figure 3: Illustration of semi-supervised contrastive learning framework for unsupervised domain adaptation. The SegCLR block corresponds to the architecture in Fig. 2. The repel losses are not used by SimSiam. Supervised losses are only used when labeled images exist. While this framework is flexible to accommodate any number of labeled images, at least one is required to drive the decoder arm of the underlying UNet.
II-D Pair generation strategies
The approach chosen for contrastive pair generation is key in self-supervised learning. We herein propose and compare different pair generation functions $P(\cdot)$ for volumetric OCT images, as illustrated in Fig. 4.
First, the augmentation-based pair formation typically employed for natural images (e.g., in SimCLR and SimSiam) is adapted for volumetric OCT images, denoted as $P_{\mathrm{a}}$. To that end, labeled slices in $D^{\mathrm{s}}$ and random slices in $D^{\mathrm{t}}$ are augmented with horizontal flipping ($p=0.5$), horizontal and vertical translation (within 25% of the image size), zoom in (up to 50%), and color distortion (brightness up to 60% and jittering up to 20%). For color augmentation, images are first transformed to RGB and then back to grayscale.
Alternatively, we propose a slice-based pairing $P_{\mathrm{s}}$ that leverages the expected coherence of nearby slices in a 3D volume for CL. For this, $x_{i}^{\prime}=x_{i}$ is chosen with a slice index $b_{i}^{\prime}$ in 3D. Then, $x_{i}^{\prime\prime}$ is a slice from the same volume with the (rounded) slice index $b_{i}^{\prime\prime}$ sampled from a Gaussian distribution $\phi$ centered on the index of the original image, i.e., $b_{i}^{\prime\prime}\sim\phi(b_{i}^{\prime},\sigma)$, with standard deviation $\sigma$ as a hyperparameter. Combining the two pairing strategies yields $P_{\mathrm{s+a}}$ where $P_{\mathrm{s}}$ is used first and the augmentations in $P_{\mathrm{a}}$ are then applied on the selected slices.
Figure 4: Proposed pair generation approaches for contrastive learning with 3D images using SegCLR.
II-E Aggregation of contrastive features
A projection head $C(\cdot)$ is formed by an aggregation function $\rho^{\mathrm{agg}}$ that aggregates features $h$ in a vector, which is then processed by a multilayer perceptron $\rho^{\mathrm{MLP}}$ to create projection $z$. Typical contrastive learning frameworks such as SimCLR and SimSiam use a projection (denoted herein by $C_{\mathrm{pool}}$) where $\rho_{\mathrm{pool}}^{\mathrm{agg}}:\mathbb{R}^{w\times h\times c}\rightarrow\mathbb{R}^{1\times 1\times c}$ is a global pooling operation on the width $w$, height $h$, and channels $c$ of the input features. Such projection $C_{\mathrm{pool}}$ may be suboptimal for learning representations to effectively leverage segmentation information, as backpropagation from $\mathcal{L}_{\mathrm{con}}$ would lose the spatial context. Instead we propose $C_{\mathrm{ch}}$, for which $\rho_{\mathrm{ch}}^{\mathrm{agg}}:\mathbb{R}^{w\times h\times c}\rightarrow\mathbb{R}^{w\times h\times 1}$ is a $1$$\times$$1$$\times$$1$ convolutional layer that learns layer aggregation while preserving spatial context.
II-F Implementation
Adam optimizer [24] was used in all models, with a learning rate of $10^{-3}$. Dropout with $p=0.5$ is applied on the layers illustrated with orange arrows in Fig. 2. The projection head $\rho^{\textrm{MLP}}$ in $C(\cdot)$ is formed by two fully-connected layer with 128 units each, where the first one uses group normalization [25] (with a group size of 4) and ReLU activation. We heuristically set $\lambda=20$ and the standard deviation for $P_{\mathrm{s}}$ as $\sigma=250\,\mu m$, which is the range for which we observe roughly similar features across slices. Supervised models were trained for 200 epochs, and the model at the epoch with the highest average Dice coefficient across classes on the validation set was selected for evaluation on a holdout test set. Our implementation is in Tensorflow 2.7, ran on an NVIDIA V100 GPU.
II-G Evaluation setup
[…]
Averaging metrics across classes with a large variation may lead to bias. Thus, for the metrics reported as boxplots, each per-slice metric $m^{c}_{i}$ for method $i$ and class $c$ is first normalized by its class Baseline $m^{c}_{\textrm{bas}}$, and then these are averaged over all $c$ and images on the test set. The average of all Baseline replicates is used as $m^{c}_{\textrm{bas}}$ for calculating the relative metrics of all other models. Absolute metrics for each class are also reported in the tables as mean $\pm$ standard deviation. To facilitate model comparison, the metrics included in most tables in the main text are averaged across classes. To further allow for the interpretation and comparison of absolute segmentation performances, we included detailed tables in Supplementary Materials with metrics separately for individual labels.
For comparing and ranking models, we adapted the strategy from [26] to simultaneously consider both Dice and UVD as complementary metrics. Accordingly, the models are first ranked among each other on individual B-scans, separately based on their Dice and UVD. These rankings are next averaged across the B-scans and the two metrics, to obtain and report a final ranking of each model for each domain.
III Dataset
We employ three large OCT datasets from clinical trials. As each dataset involve different eye diseases and/or acquisition devices, each then denotes a distinct domain $D_{i},\,i\in\{1..3\}$, with details as follows:
$\bullet$ $D_{1}$: Images of neovascular age-related macular degeneration (nAMD) patients, acquired using a Spectralis (Heidelberg Engineering) imaging device, yielding scans of $512\times 496\times 49$ or $768\times 496\times 19$ voxels, with a resolution of $10\times 4\times 111$ or $5\times 4\times 221$ $\mu$m/voxel, respectively. These were acquired as part of the phase-2 AVENUE trial (NCT02484690).
$\bullet$ $D_{2}$: Images of nAMD patients, acquired using a Cirrus HD-OCT III (Carl Zeiss Meditec) imaging device, yielding scans with $512\times 1024\times 128$ voxels and a resolution of $11.7\times 2.0\times 47.2$ $\mu$m/voxel. These were acquired as part of the phase-3 HARBOR trial (NCT00891735).
$\bullet$ $D_{3}$: Images of diabetic macular edema (DME) patients, acquired using a Spectralis device with scan sizes and resolutions the same as for $D_{1}$. These were acquired as part of the phase-2 BOULEVARD trial (NCT02699450).
All slices (B-scans) from the three datasets were resampled to $512\times 512$ pixels with the approximate resolution of $10\times 4$ $\mu$m/pixel. Selected B-scans from $D_{1}$ and $D_{2}$ were manually annotated for fluid regions of potential diagnostic value, i.e., intraretinal fluid (IRF), subretinal fluid (SRF), pigment epithelial detachment (PED), and subretinal hyperreflective material (SHRM). B-scans from $D_{3}$ were annotated for IRF and SRF, but not PED nor SHRM, as these are not relevant for diabetic macular edema patients. More details on these datasets and the annotation protocol can be found in [27].
Labeled data exists for all 3 domains as detailed in Table S1. For each experimental setting, different ablation combinations were performed by training models on labeled source domain(s) $D^{s}$ and applying the trained models on unlabeled domain(s) $D^{t}$. When a domain is considered $D^{t}$, we omit its labels in the training and use them only for ($i$) evaluating the models on the test set of that domain, or ($ii$) training of an UpperBound model used as a reference. Our data stratification is detailed in Table S1.
IV Results and Discussion
In most experimental settings, we train a model $F$ concurrently on $(x,y)\in D^{\mathrm{s}}$ and $x\in D^{\mathrm{t}}$, and then apply this trained model on the target data for evaluating $F(x\,|\,x\!\in\!D^{\mathrm{t}})\rightarrow y\in D^{\mathrm{t}}$. In the zero-shot DA setting, we train only on $(x,y)\in D^{\mathrm{s}}$, without any information from $D^{\mathrm{t}}$. We evaluate the trained models also on their initial source domain $F(x\,|\,x\!\in\!D^{\mathrm{s}})\rightarrow y\in D^{\mathrm{s}}$ to assess the retention of source-domain segmentation capability.
IV-A Joint learning for unsupervised domain adaptation across imaging devices
We first evaluate the proposed SegCLR framework in an unsupervised domain adaptation setting where the domain shift is caused by images being acquired with different imaging devices. To address the DA between Spectralis and Cirrus devices, the source and target domains were chosen as $D^{s}=D_{1}$ and $D^{t}=D_{2}$, since the latter domain has many more unlabeled images to facilitate unsupervised DA. This also helps us replicate the setting in [23] showing our preliminary results.
TABLE I: Absolute segmentation metrics averaged across classes in the unsupervised domain adaptation setting using $D^{s}=D_{1}$ and $D^{t}=D_{2}$. Results with the best (underlined) and the second best performance are in bold (excluding UpperBound).
Model Dice UVD
[…]
UpperBound 61.30$\pm$ 4.34 N/A 9.61$\pm$ 1.42 N/A
Baseline 39.23$\pm$ 8.18 67.14$\pm$ 8.24 16.47$\pm$ 2.15 6.00$\pm$ 2.32
[…]
SimSiam 41.14$\pm$ 7.86 67.40$\pm$ 8.63 15.98$\pm$ 1.89 6.03$\pm$ 2.44
[…]
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 39.39$\pm$ 8.38 52.24$\pm$ 7.35 17.25$\pm$ 6.00 8.29$\pm$ 3.27
[…]
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 50.60$\pm$ 8.98 66.38$\pm$ 7.06 13.25$\pm$ 3.70 6.13$\pm$ 2.41
[…]
A supervised UNet model, Baseline, is trained only on the source domain $D^{\mathrm{s}}$. An UpperBound supervised model trained on labeled data from $D^{t}$ is included for comparison. This labeled data is used here only as a reference and is ablated for all other models. Table I shows very poor results for Baseline on $D^{t}$, while results are much better for UpperBound on $D^{t}$ or Baseline on $D^{s}$, confirming that the two domains indeed differ from supervised learning perspective. For easier comparison, Figure 5 shows the metrics relative to the average of Baseline models and averaged across labels, i.e., so-called relative metrics. The distribution across 10 training repetitions are shown to highlight the large inherent variation in these models, including Baseline.
Learning representations of $D^{\mathrm{t}}$ with SimCLR and SimSiam with subsequent finetuning on $D^{\mathrm{s}}$ shows a slight improvement over Baseline for $D^{t}$, as seen in Figure 5 by their ($\approx$2%) higher Dice and lower UVD. This confirms that these pretraining strategies apply also on our OCT data. The results for $D^{s}$ are comparable to Baseline, as the finetuning on this domain, i.e., initializing the Baseline with the target-learned representations, does not seem to add more information than existing from the source supervision.
Figure 5: Relative segmentation metrics for models using $D^{s}=D_{1}$ and ${D^{t}=D_{2}}$. The dashed line depicts the average Baseline result used as reference for the relative metrics.
Our proposed SegCLR framework is used here for joint training, which augments SimCLR with a supervised loss and for contrasting segmentations (Section II-C). SegCLR increases the number of network parameters only for training and merely by 6.85% with respect to Baseline (UNet). The results in Figure 5 show that most choices of the projection head $P$ and the contrastive head $C$ lead to much better results (higher Dice and lower UVD) for SegCLR, notably even surpassing the UpperBound in some cases. The only configurations for which the results do not follow this positive trend involve the slice-based pairing strategy $P_{s}$ alone, indicating that merely contrasting nearby slices does not facilitate extracting features useful for segmentation. While $P_{s}$ has the potential to be a complementary augmentation to $P_{a}$, it does not seem to be useful for learning informative features by itself. This corroborates other works on augmentations [11] reporting cropping to be a crucial augmentation, which in our work is excluded when $P_{a}$ is not applied. Indeed, most SegCLR configurations yield satisfactory results, with only SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) failing substantially compared to Baseline, both for $D^{t}$ and $D^{s}$. This is likely because the pooling operation removes the spatial context between nearby and hence relatively similar slices, hence preventing the learning of relevant features from spatial correspondence. In comparison, SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) can easily capture pixel-wise differences between nearby slices, thanks to $C_{\mathrm{ch}}$ being spatially resolved, so that relevant features can be learned even with the $P_{\mathrm{s}}$ pairing.
In [23] with a similar experimental setting we reported SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$) as the superior SegCLR configuration. Our strategy herein with repetition of experiments with different initializations (10 random seeds) reveals a large dependence on initializations, where sometimes different conclusions may be reached for specific seeds. While such repeated evaluation strategy is computationally expensive, we hope this observation to serve as an incentive to standardize evaluation processes for generalizable conclusions.
Most configurations of SegCLR lead to results superior to other methods. In Table I, SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) is seen to consistently produce the best or the second best segmentation results for both $D^{s}$ and $D^{t}$ for DA across imaging devices. In particular, this configuration is seen to preserve the source domain information most successfully, with the best metric in $D^{s}$ — notably even compared to the supervised Baseline. This shows that supervised information from the labeled domain is not forgotten (e.g., as a model capacity trade-off when learning from the large unlabeled domain) and is indeed even enhanced and enriched with the added contrastive strategy. Hence, SegCLR can be used for segmentation of both $D^{s}$ and $D^{t}$, replacing the conventional Baseline approaches for the labeled domains, while approaching UpperBound such that it obviates labeling efforts for scenarios with no training data is or can be available in a target domain.
IV-B Unsupervised domain adaptation across eye diseases
This section evaluates SegCLR for domain adaptation across eye diseases, i.e., when image appearances are similar but the content differs. To that end, an experimental setting with $D^{s}=D_{1}$ and $D^{t}=D_{3}$ is used. Similarly to the previous section, Figure 6 and Table II show that the Baseline model trained on $D^{s}$ performs substantially worse on $D^{t}$ compared to an UpperBound model trained on that target domain. Interestingly, in contrast to the previous section, SimCLR, i.e., a model with contrastively pretrained features, is worse than the Baseline for both $D^{s}$ and $D^{t}$ in this setting. This can be due to instabilities in the contrastive loss, which do not affect SegCLR thanks to the proposed joint training strategy.
TABLE II: Absolute segmentation metrics averaged across classes in the unsupervised domain adaptation setting using $D^{s}=D_{1}$ and $D^{t}=D_{3}$. Results with the best (underlined) and the second best performance are in bold (excluding UpperBound).
Model Dice UVD
[…]
UpperBound 79.32$\pm$ 14.25 N/A 6.57$\pm$ 4.91 N/A
Baseline 64.77$\pm$ 11.71 67.14$\pm$ 8.24 9.38$\pm$ 7.06 6.00$\pm$ 2.32
[…]
SimSiam 65.71$\pm$ 12.23 66.90$\pm$ 7.81 9.21$\pm$ 6.87 5.95$\pm$ 2.33
[…]
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 63.13$\pm$ 15.61 63.01$\pm$ 7.41 10.49$\pm$ 7.08 6.25$\pm$ 2.40
[…]
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 65.00$\pm$ 11.09 65.85$\pm$ 7.10 9.82$\pm$ 6.68 6.20$\pm$ 2.50
[…]
Figure 6: Relative segmentation metrics for models using $D^{s}=D_{1}$ and ${D^{t}=D_{3}}$. The dashed line depicts the Baseline results used as reference for the relative metrics.
Similarly to the DA setting in the previous subsection, Figure 6 confirms that SegCLR with most of its configurations substantially outperforms Baseline for $D^{t}$, while also yielding marginally better results for $D^{s}$. Only the configurations with $P_{\mathrm{s}}$ alone leads to results inferior to Baseline, with poorest results with the SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) configuration. In particular, Table II shows SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$) overall performs the best, with SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) being a close follower. The proposed method SegCLR being successful also in this different DA setting between eye diseases also involving different label sets reinforces the conclusions for its proposed use in diverse DA settings. Indeed, its superior performance also in the source domain indicates its utility as a general learning framework for either labeled or unlabeled data from multiple domains is concerned.
IV-C Selecting the reference SegCLR implementation
To select a specific SegCLR configuration for subsequent experiments, we tabulate average rankings of the models in Table III that summarize the results of both studied DA settings from the previous two subsections.
TABLE III: Ranking (lower is better) of method variants for the experiments of domain adaptation across devices in Section IV-A ($D^{t}=D_{2}$) and across diseases in Section IV-B ($D^{t}=D_{3}$) following the reporting scheme described in Section II-G. The results are sorted by Overall ranking, which is the average ranking from both experiments across $D^{s}$ and $D^{t}$.
Model $D^{t}=D_{2}$ $D^{t}=D_{3}$ Overall
[…]
SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) 3.65 4.66 3.68 4.75 4.21
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$) 3.61 4.88 3.67 4.67 4.24
SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{pool}}$) 3.82 4.72 3.98 4.73 4.34
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{pool}}$) 3.52 4.81 4.24 4.63 4.38
SimSiam 6.17 4.37 4.89 4.71 4.91
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 4.81 4.75 5.02 5.21 4.96
Baseline 6.51 4.47 5.11 4.68 5.05
SimCLR 6.21 4.64 5.95 5.71 5.59
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 6.59 7.54 5.85 5.72 6.38
The rankings confirm the clear advantage of most SegCLR models relative to Baseline, SimCLR, and SimSiam. They also show that pretrained models do not help in such unspervised domain adaptation tasks, for which joint training with SegCLR appears to be the key (with the only exception for the use of the pairing strategy $P_{\mathrm{s}}$, as discussed above). Although the rankings differ slightly depending on whether the source or the target domain performance is prioritized, overall SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) is seen to be the best performer. By not requiring slice sampling, this model also has a simpler implementation in practice, e.g., compared to SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$), which is a top contender. In addition, the proposed $C_{\mathrm{ch}}$ adds a mere 0.03% more parameters to the baseline model. Accordingly, we chose SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) as the reference configuration for our proposed method, calling it simply SegCLR henceforth, to use in the subsequent experiments and ablations.
IV-D Effect of amount of target data on SegCLR
The success of modern self-supervised learning frameworks has been largely attributed to the availability and access to large unlabeled datasets [12, 15, 14, 13]. Previous sections demonstrate the efficacy of SegCLR given such large datasets, from extensive clinical trials which are rarely available in practice. An essential question thus remains as to how such framework performs in scenarios with smaller sets of unlabeled data. This question is studied here as an ablation study by reducing the amount of unlabeled data. We evaluate this with the experimental setting for DA across imaging devices (Section IV-A).
Figure 7 presents the results of SegCLR when trained with logarithmically reducing amount of unlabeled data compared to the earlier experiments. The experiments are taken down to the point of including no unlabaled data, i.e., the contrastive loss being computed only using the source domain data (jointly with its supervised loss). Results show that the source domain performance is not affected much by the amount of unlabeled data from a different domain, which is expected since the supervised source domain is the major source of information and the main driver for this task. Interestingly, however, regardless of the amount of unlabaled data, SegCLR performs roughly 1% in Dice on average compared to Baseline. This suggests that the proposed framework and contrastive strategy implicitly advances and improves standard UNet training and its generalization to test data from the same source domain, regardless of data existence from yet another domain.
Results in Figure 7 for the target domain of interest shows an expected trend of deterioration with diminishing amount of target data. Nevertheless, surprisingly the reduction in performance is quite minor, and a reduction of only 5.25$\pm$7.33Dice and -0.99$\pm$2.13UVD is observed between no unlabeled target data and 100% unlabeled target data (of nearly 7 million B-scans). Strikingly, SegCLR without any unlabeled data (i.e., SegCLR zero-shot learning) performs still significantly superior to Baseline, by 13.84$\pm$6.22Dice and -3.85$\pm$1.64UVD. This again suggests that the introduced framework with the joint learning strategy inherently enables better generalization, even in supervised settings without requiring additional data and without adding much to the network complexity. We further study this surprising observation in the next section.
Figure 7: Segmentation results of SegCLR for ablation experiment of unlabeled data size. The 95% confidence interval is shown as blue bands for SegCLR and as gray dotted lines for Baseline replicates. The average of Baseline replicates (black dashed line at 0) is the basis for relative comparisons, i.e., a rel. Dice above 0 or a rel. UVD below 0 indicates a superior performance by SegCLR.
IV-E Zero-shot domain adaptation on different domain combinations
In this section, we assess SegCLR zero-shot DA capability across different experimental settings. To that end, we consider all possible combinations of the datasets $D_{1}$, $D_{2}$, and $D_{3}$ as $D^{s}$ and $D^{t}$. Note that in this zero-shot setting, $D^{t}$ is exclusively reserved for evaluation, with its data (including unlabeled images) not been seen during the training phase.
Figure 8 and Table IV demonstrate the superiority of SegCLR over Baseline (UNet) consistently in nearly all experimental dataset combinations. The only setting SegCLR is clearly inferior (for both Dice and UVD) is for training on $D^{s}=D_{3}$ and evaluating on $D^{t}=D_{1}$, for which the imaging device stays the same but the disease type (labeling info and label sets) mainly differ. Note that with the imaging device DA from $D^{s}\in\{D_{1},D_{2}\}$ to $D^{t}=D_{2}$ leads a relatively larger reduction in performance for Baseline, which cannot generalize across such large domain gap — a setting in which SegCLR appears to be relatively successful. Visual examples from different experimental combinations are shown in Figure 9.
Figure 8: Results for zero-shot domain adaptation models trained on different source domains $D^{s}$ shown in the x axes, with $D^{\mathrm{All}}=D_{1}\cup D_{2}\cup D_{3}$. Each graph shows the evaluation results on each target domains $D^{t}$. The metrics are relative to the average of the Baseline models trained on the same $D^{s}$. The distributions of Baselines are also provided to illustrate the variation in Baseline repetitions. A positive Dice value for training on $D_{i}$ and evaluating on $D_{j}$ means that a zero-shot CL with SegCLR from domain $i$ applied to domain $j$ is superior to conventional supervised-only training. This indicates that the domain gap between $i$ and $j$ can be bridged by representation learning with SegCLR. Note that, although not directly observable in these normalized plots, the absolute Baseline performance for $i=j$ is much higher than otherwise, as expected. These results for $i=j$ show that SegCLR is superior to traditional supervised learning Baseline, even when applied on the trained domain itself, i.e., without any domain shift. SegCLR can thus augment and replace conventional UNet for better generalizability.
Overall, SegCLR is seen to outperform Baseline across different combinations of source and target datasets. Additionally, SegCLR surpasses Baseline even when evaluated within the same source domain. This suggests that the benefits of SegCLR extend beyond DA alone, proving valuable even for conventional supervised settings with training and deployment within the same domain, i.e., as a plug-and-play replacement for UNet.
Collectively, these results affirm the prior hypothesis that SegCLR exhibits robust performance in realistic scenarios where deployment on entirely unseen domains is anticipated, particularly when the domain gap arises from changes image apperance, e.g., across different imaging devices.
Figure 9: Qualitative assessment of SegCLR and Baseline, trained on different domains (rows) for example images from each domain (columns). Note that $D_{3}$ only has two of the four labels. These visually examplify that ($i$) training and evaluating on the same domain works well with both Baseline and SegCLR, ($ii$) SegCLR outperforms Baseline when training and evaluation domains differ, i.e., in UDA settings, ($iii$) multi-domain training $D_{\mathrm{All}}$ performs better than single-domain training, for both Baseline and SegCLR, with minor superiority of SegCLR (see bottom-right for $D_{3}$).
IV-F Multi-domain supervised contrastive learning
Although the evaluations so far above considered only single source domains, it is common in real-world contexts to have data available from multiple domains. For conventional supervised training, preliminary findings, e.g.,  [28], suggest that training with multi-domains simultaneously may yield superior segmentation outcomes on each domain separately, compared to models trained individually on each domain. This indicates that different domains may bring in complementary and supporting information in a multi-domain training setting. We test here SegCLR in such practical setting of multi-domain training.
We train both Baseline and SegCLR models on $D^{s}=D_{\mathrm{All}}\equiv D_{1}\cup D_{2}\cup D_{3}$, evaluating the two trained models on each domain separately. Note that no unlabeled data is used in this setting, where only the images from the annotated training data were used for contrastive learning. Quantitative results are shown in Figure 8 and Table IV, with examples depicted in Figure 9. The results corroborate the findings in [28] in that Baseline trained on $D_{\mathrm{All}}$ outperforms all other individual Baseline models for the corresponding target domains (i.e., , comparing the rows of Baseline across each column), except for a minuscule margin in Dice for $D_{3}$. The observation further extends to SegCLR, i.e., when trained on $D_{\mathrm{All}}$, SegCLR consistently outperforms Baseline as well as all other individual-domain-trained SegCLR results, for all examined domains and both metrics. Evidently, SegCLR framework can effectively leverage contrastive information even from labeled data, thereby enhancing model generalizability across datasets as well as in-domain image variations. Consequently, SegCLR emerges as an ideal framework also for supervised learning only with labeled data, i.e., when there is no domain shift. This further includes settings involving multiple domains for training and/or inference.
TABLE IV: Zero-shot DA results for models trained on different $D^{s}$ (rows) and evaluated on different $D^{t}$ (columns). Absolute metrics are shown averaged across all classes, in bold for the best performance.
$D^{s}$ Model Dice UVD
[…]
$D_{\textrm{1}}$ Baseline 67.14$\pm$ 8.24 39.23$\pm$ 8.18 64.77$\pm$ 11.71 6.00$\pm$ 2.32 16.47$\pm$ 2.15 9.38$\pm$ 7.06
SegCLR 69.13$\pm$ 7.94 53.07$\pm$ 8.35 71.13$\pm$ 17.16 5.94$\pm$ 2.45 12.61$\pm$ 2.24 8.27$\pm$ 6.10
[…]
SegCLR 54.55$\pm$ 6.27 64.08$\pm$ 6.23 55.11$\pm$ 22.72 8.23$\pm$ 3.60 9.12$\pm$ 1.58 11.20$\pm$ 8.26
[…]
SegCLR 40.10$\pm$ 5.25 40.50$\pm$ 9.41 80.21$\pm$ 13.02 11.00$\pm$ 3.53 25.12$\pm$ 10.40 6.06$\pm$ 4.69
[…]
V Conclusion
We have introduced SegCLR, a novel framework for learning across different domains, either labeled or unlabeled. Through a series of extensive experiments on OCT datasets from different clinical trials, including different acquisition devices and disease conditions, SegCLR with joint training is demonstrated to be superior to conventional training and contrastively-pretrained strategies. SegCLR can address significant domain shifts in unsupervised domain adaptation (UDA), as demonstrated in scenarios with changing imaging devices and eye diseases, with promising results for both source and target domains. Our experimental observations highlight the importance of training replicates in deep learning experimental work; a procedure that, while obvious, is rarely used. While the proposed SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) outperforms other configurations across all the experiments, most configurations leading to similar improvements in results indicates the robustness of the proposed framework to structural changes. Note that hyper-parameters were mostly similar across the diverse experimental settings, reaffirming the robustness.
The effectiveness of SegCLR does not rely on the availability of large amounts of unlabeled data, and it performs strikingly well even with limited or no unlabeled data from the target domain, i.e., in unsupervised and zero-shot domain adaptation scenarios. These have been demonstrated in a variety of experimental settings, with the exciting implications as follows: First, on the contrary to what can be anticipated, very large unlabeled datasets are not required to benefit from SegCLR, and even minimal target data is helpful. Second, impressive results can already be attained with no access to or knowledge of the target domain at all! In such zero-shot domain adaptation setting, labeled data is available only from source domain(s), with additional target domains only encountered upon deployment. The versatility of SegCLR is further demonstrated in a multi-domain training experiment where it is concurrently trained on labeled data from multiple domains. The results affirm that SegCLR is not only beneficial for new, unseen domains in a UDA setting, but it also yields superior results in typical supervised settings with labeled data, from both a single or multiple domains. Accordingly, we propose SegCLR to augment conventional supervised UNet, with added generalizability and without compromise.
In summary, our findings indicate that SegCLR provides a powerful and versatile approach to tackle training and inference across multiple domains in medical imaging. With its capability of learning from both labeled and unlabeled data across diverse domains, SegCLR shows great promise for improving the applicability of deep learning models to complex clinical settings involving heterogeneous domains.
Supplementary Material
TABLE S1: Datasets employed for the training and evaluation of models. Labeled data for training is displayed as #training+#validation. While the unlabeled data for $D_{2}$ and $D_{3}$ come from a database of numerous non-annotated volumes, the unlabeled data for $D_{1}$ is considered to be the unlabeled B-scans of the labeled volumes in order to enable evaluations also on that domain.
Domain Device Disease Training Testing
[…]
$D_{1}$ Spectralis nAMD 1363+243 234+41 11 466 275 163 28
$D_{2}$ Cirrus nAMD 735+125 122+21 $6.8\times 10^{6}$ 53 197 99 17
$D_{3}$ Spectralis DME 1264+226 228+40 $1.1\times 10^{5}$ 4098 196 35
TABLE S2: Absolute segmentation metrics for different tissue classes in the unsupervised domain adaptation setting with $D^{s}=D_{1}$ and $D^{t}=D_{2}$.
Model Dice UVD
[…]
UpperBound 67.25$\pm$ 1.66 59.64$\pm$ 3.31 60.39$\pm$ 1.70 57.94$\pm$ 3.02 N/A N/A N/A N/A 11.70$\pm$ 0.36 9.78$\pm$ 0.53 8.19$\pm$ 0.46 8.77$\pm$ 0.46 N/A N/A N/A N/A
Baseline 44.82$\pm$ 4.90 44.22$\pm$ 6.15 34.98$\pm$ 7.72 32.91$\pm$ 6.48 77.72$\pm$ 2.07 63.64$\pm$ 1.13 70.98$\pm$ 0.57 56.20$\pm$ 0.99 18.58$\pm$ 0.92 16.70$\pm$ 1.96 14.45$\pm$ 2.19 16.14$\pm$ 0.98 2.26$\pm$ 0.11 8.15$\pm$ 0.30 6.22$\pm$ 0.47 7.38$\pm$ 0.27
[…]
SimSiam 45.37$\pm$ 3.83 48.94$\pm$ 4.52 36.09$\pm$ 5.45 34.17$\pm$ 5.72 78.81$\pm$ 1.94 63.26$\pm$ 1.23 71.15$\pm$ 0.77 56.37$\pm$ 1.46 18.00$\pm$ 0.98 16.30$\pm$ 1.49 13.95$\pm$ 1.27 15.66$\pm$ 1.18 2.22$\pm$ 0.07 8.54$\pm$ 0.43 5.95$\pm$ 0.44 7.42$\pm$ 0.37
[…]
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 34.78$\pm$ 4.71 35.46$\pm$ 7.18 44.99$\pm$ 9.47 42.32$\pm$ 7.45 52.19$\pm$ 6.97 47.70$\pm$ 3.00 62.03$\pm$ 2.80 47.03$\pm$ 2.97 18.03$\pm$ 1.86 17.68$\pm$ 4.50 15.66$\pm$ 5.97 17.63$\pm$ 9.65 2.98$\pm$ 0.13 11.31$\pm$ 0.54 9.61$\pm$ 0.98 9.25$\pm$ 0.84
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{pool}}$) 64.00$\pm$ 6.72 59.16$\pm$ 2.88 56.50$\pm$ 3.98 58.79$\pm$ 5.43 75.51$\pm$ 3.86 62.79$\pm$ 1.70 71.30$\pm$ 1.37 57.97$\pm$ 1.18 12.49$\pm$ 0.80 10.67$\pm$ 1.49 10.99$\pm$ 2.03 12.37$\pm$ 3.20 2.63$\pm$ 0.13 8.64$\pm$ 0.49 5.56$\pm$ 0.28 7.06$\pm$ 0.42
SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) 61.29$\pm$ 7.87 58.30$\pm$ 3.06 57.09$\pm$ 4.06 56.62$\pm$ 4.57 76.22$\pm$ 5.23 63.20$\pm$ 1.84 71.31$\pm$ 2.11 61.01$\pm$ 2.12 13.07$\pm$ 1.75 10.76$\pm$ 2.12 10.96$\pm$ 2.45 11.70$\pm$ 2.79 2.52$\pm$ 0.12 8.50$\pm$ 0.58 5.40$\pm$ 0.23 6.96$\pm$ 0.41
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 52.76$\pm$ 10.43 46.87$\pm$ 10.14 51.74$\pm$ 3.14 51.03$\pm$ 10.30 74.25$\pm$ 2.87 62.46$\pm$ 0.89 71.44$\pm$ 0.82 57.36$\pm$ 1.32 14.76$\pm$ 4.17 15.02$\pm$ 4.66 11.48$\pm$ 2.55 11.74$\pm$ 1.35 2.33$\pm$ 0.08 8.50$\pm$ 0.29 6.02$\pm$ 0.31 7.66$\pm$ 0.26
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$) 62.44$\pm$ 4.73 59.30$\pm$ 6.18 55.86$\pm$ 2.05 54.78$\pm$ 6.01 75.85$\pm$ 3.82 63.79$\pm$ 1.51 69.67$\pm$ 1.96 59.86$\pm$ 2.80 13.20$\pm$ 1.81 10.75$\pm$ 1.49 11.12$\pm$ 2.30 11.17$\pm$ 1.69 2.60$\pm$ 0.17 8.81$\pm$ 0.64 5.52$\pm$ 0.39 7.37$\pm$ 0.49
TABLE S3: Absolute segmentation metrics for different tissue classes in the unsupervised domain adaptation setting with $D^{s}=D_{1}$ and $D^{t}=D_{3}$.
Model Dice UVD
[…]
UpperBound 65.45$\pm$ 1.09 93.18$\pm$ 0.85 N/A N/A N/A N/A 11.36$\pm$ 0.23 1.79$\pm$ 0.17 N/A N/A N/A N/A
Baseline 53.84$\pm$ 3.26 75.71$\pm$ 3.64 77.72$\pm$ 2.07 63.64$\pm$ 1.13 70.98$\pm$ 0.57 56.20$\pm$ 0.99 16.23$\pm$ 0.99 2.53$\pm$ 0.24 2.26$\pm$ 0.11 8.15$\pm$ 0.30 6.22$\pm$ 0.47 7.38$\pm$ 0.27
[…]
SimSiam 54.71$\pm$ 3.69 76.70$\pm$ 5.77 76.46$\pm$ 2.89 63.21$\pm$ 1.18 71.18$\pm$ 0.88 56.76$\pm$ 1.61 15.88$\pm$ 0.82 2.54$\pm$ 0.41 2.30$\pm$ 0.12 8.31$\pm$ 0.32 5.83$\pm$ 0.37 7.37$\pm$ 0.26
[…]
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 50.46$\pm$ 2.64 75.79$\pm$ 12.31 67.74$\pm$ 6.05 61.00$\pm$ 2.09 70.09$\pm$ 1.49 53.19$\pm$ 1.62 17.24$\pm$ 0.56 3.75$\pm$ 2.09 2.55$\pm$ 0.11 8.54$\pm$ 0.84 6.05$\pm$ 0.24 7.85$\pm$ 0.56
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{pool}}$) 56.45$\pm$ 4.12 86.51$\pm$ 4.18 78.03$\pm$ 2.16 64.48$\pm$ 2.08 71.06$\pm$ 1.62 60.98$\pm$ 2.91 14.12$\pm$ 0.85 2.79$\pm$ 0.71 2.50$\pm$ 0.09 8.55$\pm$ 0.84 5.33$\pm$ 0.46 6.94$\pm$ 0.55
SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) 58.15$\pm$ 3.17 88.89$\pm$ 3.21 77.82$\pm$ 2.38 63.14$\pm$ 1.86 70.92$\pm$ 2.14 61.11$\pm$ 2.04 13.59$\pm$ 0.76 2.10$\pm$ 0.54 2.51$\pm$ 0.12 8.83$\pm$ 0.61 5.23$\pm$ 0.18 7.16$\pm$ 0.47
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 56.17$\pm$ 2.47 73.84$\pm$ 8.96 73.78$\pm$ 3.06 61.66$\pm$ 2.36 70.74$\pm$ 0.95 57.21$\pm$ 2.02 16.30$\pm$ 0.69 3.33$\pm$ 0.63 2.32$\pm$ 0.08 8.81$\pm$ 0.29 6.06$\pm$ 0.54 7.62$\pm$ 0.37
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$) 58.39$\pm$ 3.46 89.60$\pm$ 3.86 79.10$\pm$ 2.48 63.72$\pm$ 1.71 70.98$\pm$ 1.84 60.88$\pm$ 2.53 13.13$\pm$ 0.62 1.99$\pm$ 0.68 2.58$\pm$ 0.20 8.32$\pm$ 0.60 5.36$\pm$ 0.59 7.10$\pm$ 0.66
TABLE S4: Zero-shot DA results for models evaluated on $D_{1}$ after training on different $D^{s}$. Absolute metrics are shown for each of the evaluated classes.
$D^{s}$ Model Dice UVD
[…]
$D_{\textrm{1}}$ Baseline 77.72$\pm$ 2.07 63.64$\pm$ 1.13 70.98$\pm$ 0.57 56.20$\pm$ 0.99 2.26$\pm$ 0.11 8.15$\pm$ 0.30 6.22$\pm$ 0.47 7.38$\pm$ 0.27
SegCLR 80.51$\pm$ 1.89 64.25$\pm$ 1.55 71.35$\pm$ 1.15 60.43$\pm$ 2.43 2.43$\pm$ 0.10 8.89$\pm$ 0.66 5.31$\pm$ 0.21 7.14$\pm$ 0.50
[…]
SegCLR 57.35$\pm$ 10.06 54.06$\pm$ 4.21 50.36$\pm$ 3.66 56.44$\pm$ 2.38 3.54$\pm$ 0.41 10.54$\pm$ 2.16 10.38$\pm$ 3.94 8.48$\pm$ 0.47
[…]
SegCLR 41.58$\pm$ 6.57 38.63$\pm$ 3.21 N/ A N/ A 8.57$\pm$ 3.56 13.42$\pm$ 0.79 N/ A N/ A
[…]
SegCLR 79.84$\pm$ 3.07 64.05$\pm$ 2.18 72.42$\pm$ 1.53 63.14$\pm$ 2.31 2.35$\pm$ 0.09 8.54$\pm$ 0.53 5.01$\pm$ 0.21 7.14$\pm$ 0.53
TABLE S5: Zero-shot DA results for models evaluated on $D_{2}$ after training on different $D^{s}$. Absolute metrics are shown for each of the evaluated classes.
$D^{s}$ Model Dice UVD
[…]
$D_{\textrm{1}}$ Baseline 44.82$\pm$ 4.90 44.22$\pm$ 6.15 34.98$\pm$ 7.72 32.91$\pm$ 6.48 18.58$\pm$ 0.92 16.70$\pm$ 1.96 14.45$\pm$ 2.19 16.14$\pm$ 0.98
SegCLR 59.66$\pm$ 7.82 52.37$\pm$ 10.61 49.93$\pm$ 3.99 50.33$\pm$ 6.67 13.77$\pm$ 1.54 12.21$\pm$ 2.92 12.54$\pm$ 2.27 11.94$\pm$ 1.86
[…]
SegCLR 71.10$\pm$ 2.44 58.19$\pm$ 5.28 66.71$\pm$ 3.26 60.34$\pm$ 2.70 10.95$\pm$ 0.54 10.03$\pm$ 0.78 7.35$\pm$ 0.38 8.17$\pm$ 0.77
[…]
SegCLR 45.37$\pm$ 6.81 35.62$\pm$ 9.37 N/ A N/ A 28.98$\pm$ 12.93 21.27$\pm$ 5.30 N/ A N/ A
[…]
SegCLR 74.27$\pm$ 1.92 67.80$\pm$ 4.46 71.05$\pm$ 2.72 64.80$\pm$ 3.22 9.58$\pm$ 0.70 8.66$\pm$ 0.38 6.98$\pm$ 0.69 7.52$\pm$ 0.37
TABLE S6: Zero-shot DA results for models evaluated on $D_{3}$ after training on different $D^{s}$. Absolute metrics are shown for each of the evaluated classes. Note that PED and SHRM are not relevant for $D_{3}$ (see Section III).
$D^{s}$ Model Dice UVD
[…]
$D_{\textrm{1}}$ Baseline 53.84$\pm$ 3.26 75.71$\pm$ 3.64 16.23$\pm$ 0.99 2.53$\pm$ 0.24
SegCLR 54.69$\pm$ 2.88 87.58$\pm$ 3.58 14.17$\pm$ 0.79 2.37$\pm$ 0.66
[…]
SegCLR 34.81$\pm$ 5.29 75.41$\pm$ 12.11 19.14$\pm$ 1.08 3.26$\pm$ 1.68
[…]
SegCLR 67.76$\pm$ 0.71 92.65$\pm$ 3.62 10.62$\pm$ 0.44 1.50$\pm$ 0.23
[…]
SegCLR 68.61$\pm$ 1.00 94.17$\pm$ 1.57 10.32$\pm$ 0.17 1.24$\pm$ 0.26